Y С B A I W310 X 38.7 a

Length AB = 8m

w = 70kN/m

a = 2.5m

E = 2.5 x 10^9 N/m^2

Determine the slopes at A & B, and vertical deflection at mid-span of the loaded W310 x38.7 beam shown using the double integration method.
Answer #1

Free Boely Dragsom Ay = Reacties force 70kN/m TT. с. 2.5m Je 5.5m A 3 at A Ayl By By – Reaction force at B loading by 70 kN/mon integrating 35(2–2:554 -35 XA + EI y = 147.65625 +632+ (2 2x3 12 12 include of at x=0 H>2.5m y=o EI (0)- 147.65625*(0) --35x83 3 VI mm2 Imid = 2.5X103 Nmm2 for slope at B put 78m Om dy ㅗ 82 dae lac=&mi -415.283203) 3 +35(8-2,5)3 277.4257302 277.